### v7

@@ Abstract
The efficient segmentation of text information from the background in degraded color document images is an important challenge in the preservation of ancient manuscripts. The imperfect preservation of ancient manuscripts has led to various types of degradation over time, such as staining, yellowing, and ink seepage, badly affecting document image binarization results. This work proposes a three-stage method to generate binarization image results for degraded colour document images using generative adversarial networks (GANs). Stage-1 involves applying discrete wavelet transform and retaining the low-low subband images for document image enhancement. In Stage-2, the original input image is split into red, green, and blue (RGB) three single-channel images and one grayscale image, and each image is trained with independent GANs to extract color foreground information. In Stage-3, the output images of Stage-2 and the resized input images are used to train independent GANs to generate document binarization results, enabling the combination of global and local features. The experimental results show that the Avg-Score of the proposed method is 77.64, 77.95, 79.05, 76.38, 75.34, and 77.00 on the (H)-DIBCO 2011, 2013, 2014, 2016, 2017, and 2018 datasets, which achieves the state-of-the-art level. The implementation code for this work is available at https://github.com/abcpp12383/ThreeStageBinarization.
@@ 1 Introduction
Historical analyses of human societies primarily rely on reconstructing and analyzing preserved ancient documents. Parchment manuscripts and paintings have predominantly served as the means of preserving crucial ancient documents throughout the evolution of human society. Analyzing preserved ancient documents presents challenges due to potential forms of interference and degradation, including paper yellowing, text fading, page contamination and moisture damage.
@@
Document image enhancement is one of the most important tasks in the reconstruction and analysis of ancient documents. It involves document image binarization to extract essential text information from the page. Elevating the image quality of deteriorated documents necessitates addressing various challenges, including text degradation and bleed-through issues, such as text degradation and bleed-through [1, 2, 3, 4]. However, traditional binarization methods cannot extract text information from the background of degraded documents effectively, often resulting in incomplete removal of shadows and noises, or loss of text information [5, 6, 7].
With the development of deep learning, neural networks have been applied to image binarization. Calvo-Zaragoza and Gallego [8] introduced a selectional auto-encoder approach, parsing the text through the trained network model, and performing binarization using threshold. Fully convolutional network (FCN) [9] enables the generated images to maintain the same size as the input images, and the task [10] of generating document binarization using FCN is also included in the semantic segmentation task. However, most image binarization methods primarily process grayscale documents, targeting contaminated black and white scanned ancient documents as their default.
@@
Considering that the images of some ancient documents are in colour, we propose a novel three-stage method for document image enhancement and colour degradation document binarization. The proposed method consists of three stages. In Stage-1, we split the original input image into red, green, and blue (RGB) three single-channel images and one grayscale image, and apply discrete wavelet transform (DWT) to retain the low-low (LL) subband images and normalize them to achieve image enhancement. In Stage-2, we train four independent adversarial networks with the four single-channel images to eliminate background information and extract color foreground details from the local image patches. During the processing of Stage-3, the multiscale adversarial networks produce $224\times 224$ image patch (local) prediction outputs and $512\times 512$ resized image (global) prediction outputs for the document images, which are subsequently fused to obtain the final output.
@@
To evaluate the performance of the proposed method for degraded document image binarization, we compare it with other methods using DIBCO datasets of 2011, 2013, 2014, 2016, 2017, and 2018 [11, 12, 13, 14, 15, 16].
@@
  • Introducing a novel three-stage generative adversarial networks method for enhancing and binarizing the degraded color documents.
  • The incorporation of DWT prior to the training of generative adversarial networks makes the image to be input closer to the Ground Truth image.
  • The proposed method achieves the state-of-the-art performance on multiple benchmark datasets.
The rest of this paper is organized as follows: Section 2 describes the deep learning method for document image binarization. Section 3 introduces a novel three-stage generative adversarial networks method. Section 4 presents the experiments performed on this work, and compares the performance of our method and other methods on several benchmark datasets. Finally, Section 5 discusses the conclusion and the future work of this paper.
@@ 2 Related Work
In previous researches focusing on document image binarization, the field is broadly divided into two categories: traditional image processing methods and deep learning-based semantic segmentation methods. However, traditional image processing binarization methods are always based on the pixel-level local threshold values, such as global binarization [6] and local adaptive binarization [5, 7]. These methods are not satisfactory when dealing with documents affected by excessive interference or degradation [17, 18].
@@
With the advancement of deep learning, the convolutional neural network (CNN) framework [19, 20] has become the predominant backbone of computer vision tasks. Within computer vision, semantic segmentation is one of the most important research topics, encompassing document image binarization. For instance, Tensmeyer et al. [10] employed pseudo-F-measure (p-FM) and F-measure (FM) losses during the training of FCN to apply the model to the document image binarization task. Ronneberger et al. [21] extended FCN, presenting the U-Net architecture, which captures contextual and spatial information through its distinctive U-shaped design. Inspired by this, Peng et al. [22] introduced an encoder-decoder architecture for document image binarization, where the decoder transforms the low-resolution representation to the original dimensions, resulting in the binarized image. Vo et al. [23] first applied hierarchical deep supervised network (DSN) to document image binarization and achieved the state-of-the-art (SOTA) performance on several benchmark datasets. He et al. [24] proposed an iterative deep learning framework applied to document image binarization and the trained model achieved excellent performance.
@@
In recent years, generative adversarial networks (GANs) [25] have achieved exciting results in the generation of real-world images. Unlike GANs which generate output images from random noise, Isola et al. [26] proposed Pix2Pix GAN based on conditional GAN (cGAN) [27] for image-to-image translation. To ensure that the generated images are more realistic, Pix2Pix GAN employs generators for loss function optimization. Within the network architecture, the generator evaluates the minimized generated image against the target output image, quantifying the $L1$ loss between them. Meanwhile, the discriminator assesses the credibility of the generated image, relying on the input image and the reference target image. This process is depicted by the subsequent equation:
$\begin{split}\emph{arg}\min_{G}\max_{D}V(G,D)=\mathop{\mathbb{E}_{x}}[\log(1-D(x,G(x)))]\\ +\mathop{\mathbb{E}_{x,y}}[\log D(x,y)]+\lambda\mathop{\mathbb{E}_{x,y}}[||y-G(x)||_{1}]\end{split}$ (1)
@@
where the input data distribution $\mathop{\mathbb{P}_{x}}$ samples the input image $x$; the ground truth image of the input image is $y$, and the hyperparameter to increase the model regularization effect is $\lambda$. Generator $G$ generates the input image as the generated image $G(x)$. From Eq. (1), the loss function between the generated image and the ground truth image is $L1$ instead of $L2$, which is beneficial to reduce blurring and ambiguity in the generation process.
@@
Bhunia et al. [28] performed document image enhancement on the DIBCO dataset and used cGAN [27] for image binarization. Inspired by Pix2Pix GAN [26], Zhao et al. [29] proposed a cascade generator structure for document image binarization based on Pix2Pix GAN to solve the problem of multi-scale information combination. De et al. [30] designed two discriminators to combine local and global information. Suh et al. [31] set up a two-stage generative adversarial network for image binarization using Patch GAN [26]. These methods keep breaking the SOTA performance on DIBCO datasets.
Figure 1: The architecture of Stage-1 of the proposed method. The original input image is divided into patches of $224\times 224$ size, and split into grayscale images and RGB three single-channel images. The single-channel images are all in grayscale. However, to visually represent the different channels of the image separately, the single-channel images in this diagram are replaced with images in three different RGB colors.
@@
Figure 2: The architecture of the Haar wavelet transform used by the proposed method, which takes the red single-channel image patch as an example. The processes of the transform of the green and blue single-channel image patches are the same.
@@ 3.1 Overall Network Architecture
Considering the complexity of color document image degradation and the advantages [31] of processing different color channels separately, our proposed method divides the overall network architecture into three stages, which focus on enhancing image quality of degraded color documents and extracting text information from the images. In Stage-1, we use DWT and normalization to preprocess the patches of the input image. These preprocessed image patches are employed to train as the input of the generator within the adversarial network framework. Stage-2 focuses on the elimination of background color noise and the extraction of color foreground information, while Stage-3 involves the fusion of local and global predictions to produce the final output. The detailed introduction of each stage can be found in Sections 3.2, 3.3, and 3.4, respectively.
@@ 3.2 Stage-1
This work employs four different generators to extract foreground color information from degraded document images while eliminating the background color. To preprocess the input of the different generators, we first split the original color image (24-bit) into RGB three single-channel images (8-bit) and one grayscale image (8-bit).
The ground truth (GT) image provided by the dataset is a single-channel binary image (8-bit), where pixel points have values of 0 for characters and 1 for the background. Images with varying channel numbers cannot be directly put into the discriminator for processing. As shown in Figure 1, we first split the RGB three channels of the original input image to create three single-channel RGB images. Then we apply DWT to each single-channel image, and retain the LL subband images for normalization. The GT images and the preprocessed grayscale images are summed at pixel wise as the corresponding ground truth images. Since the sizes of the retained LL subband images are $112\times 112$ (half of the $224\times 224$ length and width), we resize the images to the designated size of $224\times 224$ using bicubic interpolation [32].
@@ 3.2.1 Discrete Wavelet Transform
In this work, we use Haar wavelet transform [33], which is one of the DWT, to process the input image. This process can be conceived as a bandpass filter, selectively allowing signals with frequencies akin to those of the wavelet basis function to traverse. The flowchart of this DWT processing is illustrated in Figure 2, which generates four subband images: one low-frequency subband image and three high-frequency subband images. This work takes the single-channel split images as input and processes the n-direction by low-pass filtering, as shown in the following equation:
@@
Next, this work uses the low-pass filtering applied to the m-direction of Eq. (2) to obtain the following equation:
@@
In contrast to the ground truth images provided by the dataset, the RGB single-channel images split from the original input images exhibit some substantial noise. This noise prevents them from being used directly in generating corresponding ground truth images through the generator and could potentially disrupt the performance of the trained adversarial network. DWT decomposes the image into both low-frequency and high-frequency information. The low-frequency component represents the mean value and the high-frequency component represents the difference value. The mean value, classified as low-frequency information, captures gradual changes and encodes both the picture’s contour information and approximate details. Conversely, the difference value, characterized as high-frequency information, captures rapid changes and encodes intricate picture details as well as local information (including noises). Therefore, we apply DWT and retain the LL subband images to input Stage-2, as shown in Figure 2, which effectively preserves contour information and reduces noise interferences.
Figure 3: Results of Haar wavelet transform & normalization of RGB single-channel images split from the example image patches.
Figure 4: Image normalization from image HW15 of DIBCO 2011 dataset, the above is an example of no normalization after DWT retains LL subband images, whereas the below is the example of normalization after DWT retains LL subband images.
@@ 3.2.2 Normalization
Normalization entails constraining data within a desired range, with the goal of simplifying the data processing and expediting convergence. The proposed equation for this process is expressed as:
@@
In the proposed method, we normalize the LL subband images to present initially incomparable data for comparison. The reason we choose the normalized LL subband image as the output is shown in Figure 3, where we apply DWT and normalization to each subband images.
The illustrative example of the impact of normalization on RGB single-channel images after the DWT operation is provided in Figure 4, when applied to image HW15 from the DIBCO 2011 dataset, normalization effectively balances the RGB. In contrast, without applying normalization results in an output image, it heavily skews towards the blue component. In addition, we also show that the global binarization results of the images using normalization are more similar to the Ground Truth images provided by the dataset compared to the images only processed by DWT, and the experiments are explained in detail in Section 4.5.
Figure 5: The architecture of Stage-2 of the proposed method. This method employs the U-Net++ [34, 35] architecture with EfficientNet-B6 [36] as the generator and improved PatchGAN [26] as the discriminator. The ground truth image during the training process is obtained from pixel-wise summation of the dataset provided image and the output from Stage-1.
Figure 6: The architecture of Stage-3 of the proposed method. This stage comprises both local and global predictions. The local prediction involves training several $224\times 224$ image patches, while the global prediction involves directly resizing the original image to $512\times 512$ for training.

### v8

@@ Abstract
The efficient extraction of text information from the background in degraded color document images is an important challenge in the preservation of ancient manuscripts. The imperfect preservation of ancient manuscripts has led to different types of degradation over time, such as page yellowing, staining, and ink bleeding, seriously affecting the results of document image binarization. This work proposes an effective three-stage network method to image enhancement and binarization of degraded documents using generative adversarial networks (GANs). Specifically, in Stage-1, we first split the input images into multiple patches, and then split these patches into four single-channel patch images (gray, red, green, and blue). Then, three single-channel patch images (red, green, and blue) are processed by the discrete wavelet transform (DWT) with normalization. In Stage-2, we use four independent generators to separately train GAN models based on the four channels on the processed patch images to extract color foreground information. Finally, in Stage-3, we train two independent GAN models on the outputs of Stage-2 and the resized original input images ($512\times 512$) as the local and global predictions to obtain the final outputs. The experimental results show that the Avg-Score metrics of the proposed method are 77.64, 77.95, 79.05, 76.38, 75.34, and 77.00 on the (H)-DIBCO 2011, 2013, 2014, 2016, 2017, and 2018 datasets, which are at the state-of-the-art level. The implementation code for this work is available at https://github.com/abcpp12383/ThreeStageBinarization.
@@ 1 Introduction
Historical analysis of ancient human society mainly relies on the reconstruction and analysis of preserved ancient documents, where the parchment manuscripts are an important tool for preserving important ancient documents. Analyzing preserved ancient documents, especially the parchment manuscripts, presents significant challenges due to potential forms of degradation, including paper yellowing, text fading, page contamination, and moisture damage [1, 2].
@@
Document image enhancement and binarization are important steps in analyzing ancient documents, with the purpose of extracting useful text information from the background. However, improving the image quality of degraded documents requires solving several problems, such as text degradation and ink bleeding [3, 4]. Traditional threshold methods cannot effectively extract text information from the background of degraded documents, often resulting in incomplete removal of shadows and noise, or loss of text information [5, 6, 7].
With the development of deep learning, neural networks have been applied to image binarization. The common method is to apply neural networks to analyze the document images, for instance, Calvo-Zaragoza and Gallego [8] introduce a selectional auto-encoder approach to analyze document images through the trained network model, and binarize the document images using the threshold.
The classic network in semantic segmentation tasks, fully convolutional network (FCN) [9], can also be used to segment foreground and background information in document images. Tensmeyer et al.’s [10] research also shows that using FCN for document binarization can achieve good results. However, most current methods mainly process gray-scale document images, targeting contaminated black and white scanned ancient documents as their default.
@@
Considering some ancient document images are in color, we propose a novel three-stage network method for both gray-scale and color degraded document image enhancement and binarization. Specifically, the proposed method consists of three stages. In Stage-1, we first split the input images into multiple patches, and then split these patches into four single-channel patch images (gray, red, green, and blue). After that, we apply the discrete wavelet transform (DWT) on three single-channel patch images (red, green, and blue) to retain the low-low (LL) subband images, and then normalize them. To remove the noise from the background of the original input images, and to extract the foreground information from the patch images, in Stage-2, we train the GANs model on four single-channel images by using four independent generators to generate the enhanced images. In addition, in Stage-3, we use the entire document images composed of the patch images (output of Stage-2) as the local predictions, and the $512\times 512$ resized input images as the global predictions to train two independent GAN models, respectively. The outputs of these two GAN models are fused to obtain the final output images.
@@
To evaluate the performance of the proposed method for document image binarization, we compare it with other methods using the (H)-DIBCO datasets of 2011, 2013, 2014, 2016, 2017, and 2018 [11, 12, 13, 14, 15, 16].
@@
  • Introduces an effective three-stage network method that effectively performs enhancement and binarization on gray-scale and color degraded document images.
  • Shows that, compared with other methods using the same training set, the proposed method achieves the state-of-the-art performance on multiple benchmark datasets.
  • Demonstrates that for document images, the binarization results of the images processed by the discrete wavelet transform with normalization are closer to the Ground-Truth images compared to the binarization results of the original input images.
The rest of this paper is organized as follows: Section 2 describes the deep learning methods for document image enhancement and binarization. Section 3 details the proposed three-stage network method. Section 4 presents the results of the image enhancement experiment and the comparison experiment, and shows the comparative results of the performance of the proposed method and other methods on multiple benchmark datasets. Finally, Section 5 discusses the conclusion and future work of this work.
@@ 2 Related Work
Document image binarization methods can be broadly divided into two categories: traditional threshold methods and deep learning methods. Among them, traditional threshold methods are mainly based on calculating the threshold value at the pixel-level of the image, such as global binarization [6] and local adaptive binarization [5, 7]. However, these methods have unsatisfactory results for documents affected by excessive interference or degradation [17, 18].
@@
Neural networks [19, 20] continue to achieve excellent results in computer vision tasks. Among them, neural networks used in semantic segmentation tasks are also applied to document image binarization. For instance, Tensmeyer et al. [10] apply FCN [9] to document image binarization, and combine pseudo-F-measure (p-FM) and F-measure (FM) losses to train the model. Inspired by U-Net [21], Peng et al. [22] introduce an encoder-decoder architecture for document image binarization, where the decoder transforms the low-resolution representation to the original dimensions, resulting in the binarized images. Vo et al. [23] apply hierarchical deep supervised network (DSN) to document image binarization, which achieves the state-of-the-art (SOTA) performance on multiple benchmark datasets. He et al. [24] propose an iterative deep learning framework applied to document image binarization, which also achieves the excellent performance.
@@
In recent years, GANs [25] have achieved excellent results in image generation tasks. Different from generating output images from random noise, Isola et al. [26] propose Pix2Pix GAN based on conditional generative adversarial network (cGAN) [27] for image-to-image translation. To ensure that the generated images are more realistic, Pix2Pix GAN employs generators to optimize the loss function. In the network architecture, the generator evaluates the minimization of generated images against the target output images, quantifying the $L1$ loss between them. Meanwhile, the discriminator evaluates the credibility of the generated images based on the input images and the reference target images (the Ground-Truth images). The process is described by the following equation:
$\begin{split}\emph{arg}\min_{G}\max_{D}V(G,D)=\mathop{\mathbb{E}_{x}}[\log(1-D(x,G(x)))]\\ +\mathop{\mathbb{E}_{x,y}}[\log D(x,y)]+\lambda\mathop{\mathbb{E}_{x,y}}[||y-G(x)||_{1}],\end{split}$ (1)
@@
where the input data distribution $\mathop{\mathbb{P}_{x}}$ samples the input images $x$, and the Ground-Truth images are $y$. The hyperparameter to increase the model regularization effect is $\lambda$. Generator ($G$) generates the input images as the generated images $G(x)$. From Eq. (1), the loss function between the generated images and the Ground-Truth images is $L1$ instead of $L2$, which is beneficial to reduce blurring and ambiguity in the generation process.
@@
Therefore, in addition to neural networks applied in semantic segmentation tasks, GANs have also achieved satisfactory results for document image binarization. Bhunia et al. [28] employ cGAN [27] on the DIBCO datasets for document image enhancement and binarization. Inspired by Pix2Pix GAN [26], Zhao et al. [29] propose a cascade generator structure for document image binarization based on Pix2Pix GAN to solve the problem of multi-scale information combination. De et al. [30] design two discriminators to combine local and global information. Suh et al. [31] propose a novel two-stage generative adversarial network method using PatchGANs [26] for document image enhancement and binarization. Souibgui et al. [32] introduce an efficient end-to-end framework named document enhancement generative adversarial networks (DE-GAN). This framework is designed to recover severely degraded document images by generating high-quality enhanced versions of these documents. These methods keep breaking the SOTA performance on the DIBCO datasets.
In addition, diffusion models have also begun to be applied to document image binarization. For instance, Cicchetti and Comminiello [33] propose a novel generative framework based on the diffusion probabilistic model (DPM) for document image enhancement, which includes an efficient nonlinear activation free (NAF) network and an additional differentiable module based on the convolutional recurrent neural network (CRNN).
Figure 1: The overall three-stage network architecture of the proposed method for document image enhancement and binarization.
Figure 2: The architecture of Stage-1 of the proposed method. The original input images are split into multiple patches ($224\times 224$), and split into four single-channel patch images (gray, red, green, and blue), where three single-channel patch images (red, green, and blue) are processed by the Haar wavelet transform with normalization.
@@
Figure 3: The flow chart of the Haar wavelet transform with normalization used by the proposed method, which takes the red single-channel patch images as the example. The processes of the green and blue single-channel patch images are the same.
Figure 4: Example of the Haar wavelet transform with normalization of three single-channel patch images (red, green, and blue).
Figure 5: Example of the ablation study on image normalization, the above figure presents the image only processed by the Haar wavelet transform, and the below figure presents the image processed by the the Haar wavelet transform with normalization.
Figure 6: The architecture of Stage-2 of the proposed method. This method employs the U-Net++ [34, 35] architecture with EfficientNet-B6 [36] as the generator, and the improved PatchGAN [26] as the discriminator. The Ground-Truth images for three single-channel patch images (red, green, and blue) are the binarization results of the outputs from Stage-1.
Figure 7: The architecture of Stage-3 of the proposed method. This stage comprises the local prediction (left-side) and the global prediction (right-side).
@@ 3.1 Overall Network Architecture
Considering the complexity of degradation of color documents, better results are often obtained by processing images of different color channels separately [31]. Therefore, this method divides the overall network architecture into three stages, which are used for document image enhancement, text information extraction, and document image binarization, as shown in Figure 1. The images processed in Stage-1 are prepared for training GANs models in Stage-2 and Stage-3. Stage-2 focuses on the removal of noise from the background and the extraction color foreground information, while Stage-3 involves the fusion of the local and global predictions to obtain the final output images. The detailed introduction of each stage can be found in Sections 3.23.3, and 3.4, respectively.
@@ 3.2 Stage-1
This work employs four independent generators to train GANs model, which extracts foreground information from degraded document images while removing noise from the background. For color degraded documents, to process the inputs of different independent generators, we first split the original input images (24-bit) into four single-channel patch images (gray, red, green, and blue), each of which is 8-bit. However, for the gray-scale (8-bit) degraded documents, we only process the gray single-channel images in Stage-1 and train the model on gray single-channel images for Stage-2.
Next, we will focus on the case of color degraded documents. Since the Ground-Truth images provided by the dataset are binarized images (8-bit), images with different channel numbers cannot be directly put into the discriminator for processing. As shown in Figure 2, except for the gray single-channel images, we process the remaining three single-channel images (red, green, and blue) using the Haar wavelet transform (one of the DWTs), and retain the LL subband images for normalization. Since the sizes of the retained LL subband images are $112\times 112$ (half of $224\times 224$), we resize the images to the original size of $224\times 224$ using the bicubic interpolation [37].
In addition, as illustrated in Figure 2, the Ground-Truth images used for the four channels are different. For the gray channel, we directly use the Ground-Truth images provided by the dataset. For the remaining three channels (red, green, and blue), we employ the results of global binarization of the single-channel images processed by the DWT with normalization as the Ground-Truth images. The reasons for using DWT with normalization and their effects on the images are described in Section 4.5.
@@ 3.2.1 Discrete Wavelet Transform
In this work, we use the Haar wavelet transform [38] (one of the DWTs) to process the input images. This processing can be equated to using a bandpass filter that selectively allows signals with frequencies similar to the wavelet basis functions to pass through. The flowchart of this processing is illustrated in Figure 3, which generates four subband images, such as the low-low (LL) subband images, the high-low (HL) subband images, the low-high (LH) subband images, and the high-high (HH) subband images. The low-pass filtered signal $V_{L}$ at position $(m,n)$ is obtained by convolving the input signal $x$ with the low-pass filter $g$, where the input signal is downsampled by a factor of 2, as shown in the following equation:
@@
where $x[m,2n-k]$ is the input signal $x$ at position $(m,2n-k)$, and this indexing suggests downsampling by a factor of 2. $g[k]$ is the filter coefficients for the low-pass filter, and $K$ is the length of the filter $g$. The low-pass filtered signal $V_{L}L$ at position $(m,n)$ is obtained by convolving the intermediate low-pass filtered signal $V_{L}$ (filtered along the columns) with the low-pass filter $g$, where the intermediate signal is downsampled by a factor of 2 along the row dimension, as shown in the following equation:
@@
In contrast to the Ground-Truth images provided by the dataset, the three single-channel images (red, green, and blue) possess a significant amount of noise. The noise will prevent them from being used directly through the generator to generate the corresponding Ground-Truth images and possibly severely affect the performance of the trained GANs models. The DWT decomposes the images into both low-frequency and high-frequency information, where the low-frequency information represents the mean value, and the high-frequency information represents the difference value. The mean value captures gradients and encodes both contour information and proximate details of the images, and the difference value captures rapid changes and encodes complex image details and local information. Therefore, we apply the DWT and retain the LL subband images, as shown in Figure 3, which effectively preserves contour information and reduces the noise interference.
@@ 3.2.2 Normalization
Normalization entails constraining the data to the desired range with the aim of simplifying data processing and speeding up convergence. The proposed equation for this process is expressed as:
@@
In our method, we normalize the LL subband images by the DWT to present initially incomparable data for comparison. The reason we choose the normalized LL subband images as the output of Stage-1 is shown in Figure 4, where the foreground information in another subband images is not prominent compared to the LL subband images.
Figure 5 illustrates the example of the effect of normalization on three single-channel patch images (red, green, and blue) processed by the DWT, which effectively balances three colors when applied to the “HW15” image of the DIBCO 2011 dataset [11]. In contrast, the output images without normalization are heavily biased towards the blue component. In conclusion, the main difference between the results of Figure 5 (a) and (b) is that the images without normalization may lead to noise in the single-channel images, which seriously affect the pixel-level summation results.
In addition, we also show that, compared to the binarization results of the images only processed by the DWT, the binarization results of the images using the DWT with normalization are more similar to the Ground-Truth images provided by the dataset. The related experiments are explained in detail in Section 4.5.